OpenStudy (anonymous):

Simplify. (2+i)(2-i)(4+3i)

OpenStudy (anonymous):

\[20+15i\]

OpenStudy (anonymous):

5(4+3i) =20+15i

OpenStudy (anonymous):

start with the fact that \[(a+bi)(a-bi)=a^2+b^2\] so rewrite this as \[5(4+3i)\]

OpenStudy (anonymous):

(2+i)(2-i)=2^2-i^2... but after that... quite clueless... is this how to start off?

OpenStudy (amistre64):

i gotta wonder what constitutes a simplification ...